OpenStudy (anonymous):

how would you work this partial fraction problem? sqrt(2y^2-3)/y^2

OpenStudy (amistre64):

partial decomp you mean?

OpenStudy (amistre64):

\[\frac{\sqrt{2y^2-3}}{y^2}\] ???

OpenStudy (anonymous):

your question is not complete.

OpenStudy (anonymous):

yes that but with the dy as a variable and is a normal intergration

OpenStudy (anonymous):

as in you put +C at the end

OpenStudy (amistre64):

then this is the question \[\int\frac{\sqrt{2y^2-3}}{y^2}dy\]

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

\[\frac{\sqrt{2y^2-3}}{y^2}=\frac{1}{y^2}\sqrt{2y^2-3}\] \[\frac{1}{y^2}\sqrt{2y^2-3}=\sqrt{\frac{1}{y^4}(2y^2-3)}\]

OpenStudy (amistre64):

maybe as a start :)

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

but starts arent always good for me :) http://www.wolframalpha.com/input/?i=int+sqrt%282y^2-3%29%2Fy^2+dy looks to be a trig substitution

OpenStudy (amistre64):

it tends to get messier from there ...

OpenStudy (anonymous):

Try y = sin theta.....

OpenStudy (anonymous):

Seems to work out quite nicely with a bit of jiggling.